# Non-toric brane webs, Calabi–Yau 3-folds, and 5d SCFTs

### v1

@@ 6.2. Calabi–Yau 3-folds as non-toric deformations of toric Calabi–Yau 3-folds
in $\mathbb{C}^{3}$ by the group $\mathbb{Z}/n\mathbb{Z}$ acting with weights $(1,-1,a)$ on $\mathbb{C}^{3}$. It follows from this description that $(\overline{X}_{\mathcal{C}},\partial{\overline{X}}_{\mathcal{C}})$ admits a natural one-parameter deformation given by quotienting the hypersurface
@@
by the action of the group $\mathbb{Z}/n\mathbb{Z}$ acting with weights $(1,-1,a)$ on $\mathbb{C}^{3}$. For $n=1$, this is just smoothing the normal crossing divisor $xy=0$ into $xy=t$. If $n>1$, then, both the singular toric surface $\overline{X}_{\mathcal{C}}$ and its toric boundary divisor $\partial\overline{X}_{\mathcal{C}}$ become smooth for $t\neq 0$. According to [6, Lemma 4.1], there is no obstruction to apply the previous deformation to all the neighborhoods $(\overline{X}_{\mathcal{C}},\partial{\overline{X}}_{\mathcal{C}})$ of the 0-dimensional strata $x_{\mathcal{C}}$ to obtain a global deformation of the toric surface $(\overline{X}_{0}^{v},\partial\overline{X}_{0}^{v})$, and we denote by $(X_{0}^{v},\partial X_{0}^{v})$ the resulting log Calabi–Yau surface. This surface is smooth since all the possibly singular 0-dimensional strata of $(\overline{X}_{0}^{v},\partial\overline{X}_{0}^{v})$ become smooth in the deformation. Here, we are using the terminology of “log Calabi–Yau” surface in a sense slightly more general than in §3.1.2: the surface $(\overline{X}_{0}^{v},\partial\overline{X}_{0}^{v})$ can be non-compact, which is the case if $v$ is on the boundary of $P$, and the boundary $\partial\overline{X}_{0}^{v}$ can be smooth, which is the case if $L^{2}=0$ and all the cones in the star of $v$ are cut out.
In the web picture, the vertex $v$ of $\mathcal{P}$ corresponds to a face of the polyhedral decomposition defined by $W^{\prime}$. Moreover, this face is obtained from a face of the polyhedral decomposition defined by $W$ by “corner smoothings” induced by pushing-in 7-branes as in Figure 6.2.
[…]
Figure 6.2. The log Calabi–Yau surface $(X_{0}^{v},\partial X_{0}^{v})$ from “smoothing corners” of the toric surface $(\overline{X}_{0}^{v},\partial\overline{X}_{0}^{v})$.
The log Calabi–Yau surface $(X_{0}^{v},\partial X_{0}^{v})$ can be described explicitly as follows. The singularity of the integral affine manifold $P$ at $v$ can be thought of as several focus-focus singularities brought together, one for each cut triangle with tip at $v$. Moving slightly away from $v$ each of these focus-focus singularities along their monodromy invariant directions, the vertex $v$ becomes a smooth point of the integral affine structure, and so the star of $\mathcal{P}$ at $v$ can be viewed as the fan of a toric variety. Then, $(X_{0}^{v},\partial X_{0}^{v})$ is obtained from this toric variety as follows:
@@
    First, blow up torically so that all the half-lines starting at $v$ and passing through a focus-focus singularity are part of the fan.
@@
    Then, blow up non-torically for each focus-focus singularity a point on the component of the toric boundary corresponding to the half-line passing through the focus-focus singularity
@@
    Finally, contract all the components of the boundary that do not correspond to edges of $\mathcal{P}$.
In general, there exist several ways to split the integral affine singularity at $v$ into focus-focus singularities, which correspond to different toric models of the log Calabi–Yau surface $(X_{0}^{v},\partial X_{0}^{v})$.
Step II: Construction of the normal crossing surface $\mathcal{X}_{0}$. By Step I, we have attached a smooth log Calabi–Yau surface $(X_{0}^{v},\partial X_{0}^{v})$ to every vertex of $\mathcal{P}$, with dual intersection given by the star of $\mathcal{P}$ at $v$. In particular, the edges of $\mathcal{P}$ adjacent to $v$ correspond to irreducible components of the boundary $\partial X_{0}^{v}$. Gluing the log Calabi–Yau surfaces $(X_{0}^{v},\partial X_{0}^{v})$ along the components of their boundaries corresponding to the same edges of $\mathcal{P}$, we obtain a singular surface $\mathcal{X}_{0}$, with irreducible components $(X_{0}^{v},\partial X_{0}^{v})$ and intersection complex $\mathcal{P}$. Equivalently, the intersection complex of $\mathcal{X}_{0}$ is the polyhedral decomposition determined by the web $W^{\prime}$. Since the faces of $\mathcal{P}$ are triangles (the web $W^{\prime}$ is 3-valent), there are at most triple intersections between the irreducible components of $(X_{0}^{v},\partial X_{0}^{v})$. Since these irreducible components are smooth, we conclude that $\mathcal{X}_{0}$ is a normal crossing surface.
Step III: Construction of the Calabi–Yau 3-fold $\mathcal{X}$. We obtain a smooth non-compact Calabi–Yau 3-fold $\mathcal{X}$ as the total space of a one-parameter smoothing of the normal crossing surface. Let $(Y,D,L)$ be the log Calabi–Yau surface associated to the asymptotic web of 5-branes with 7-branes $W^{\mathrm{asym}}$. According to [36] if $L\cdot D=0$ and [6, Theorem 3.36] if $L\cdot D>0$, there exists a smooth non-compact Calabi–Yau 3-fold $\mathcal{X}$ and a holomorphic map
@@
where $\Delta$ is an open disk in $\mathbb{C}$ containing $0\in\mathbb{C}$, such that the central fiber $\pi^{-1}(0)$ is a reduced normal crossing divisor isomorphic to $\mathcal{X}_{0}$. Moreover, it is proved in [37, Proposition 3.14] if $L\cdot D=0$ and [7] if $L\cdot D>0$ that, for $t\neq 0$, the general fiber $\pi^{-1}(t)$ is diffeomorphic to $U=Y\setminus D$. Finally, by [6, Theorem 3.29], there exists a unique holomorphic map
@@
such that $\pi=\pi^{\mathrm{can}}\circ f$. Moreover, $\mathcal{X}^{\mathrm{can}}$ is non-compact Calabi–Yau 3-fold with Gorenstein canonical singularities, and $f:\mathcal{X}\rightarrow\mathcal{X}^{\mathrm{can}}$ is a crepant resolution. By construction, $\mathcal{X}^{\mathrm{can}}$ is a generally non-toric deformation of the toric 3-fold singularity $\overline{\mathcal{X}}^{\mathrm{can}}$ associated as in §2.3 to the asymptotic web of 5-branes $\overline{W}^{\mathrm{can}}$, and the crepant resolution $f:\mathcal{X}\rightarrow\overline{X}^{\mathrm{can}}$ is a deformation of the crepant toric partial resolution $\overline{\mathcal{X}}\rightarrow\overline{\mathcal{X}}^{\mathrm{can}}$ associated to the web of 5-branes $\overline{W}$. The toric Calabi–Yau 3-fold $\overline{\mathcal{X}}$ is singular in general since $\overline{W}$ might contain 3-valent vertices of multiplicity $>1$ or 4-valent vertices, and so $\mathcal{X}$ is a smoothing of $\overline{\mathcal{X}}$.
@@
Assume that $L^{2}=0$, that is, as reviewed above Equation (4.10) that there exists an elliptic fibration $c:Y\rightarrow\mathbb{P}^{1}$ with $D=c^{-1}(\infty)$ and $L=\mathcal{O}(kD)$. In particular a general curve $C\in|L|$ consists of $k$ disjoint elliptic fibers. In this case, the web $W^{\prime}$ obtained after pushing fully the 7-branes does not contain any vertex, and is topologically the union of $k$ concentric circles, which can be viewed as the tropicalization of the $k$ elliptic connected components of $C$. If $k=1$, then the complement of $W^{\prime}$ contains a unique bounded region, corresponding to a compact component of $\mathcal{X}_{0}$, which is necessarily a del Pezzo surface $S$. It follows that $\mathcal{X}$ is the total space $K_{S}$ of the canonical divisor of $S$, that is, a “local del Pezzos surface”, and the map $\pi:\mathcal{X}\rightarrow\Delta$ is induced by a general anticanonical section, whose zero-locus is a smooth genus one curve $E$. If $(Y,D,L)$ corresponds to the $E_{Q-3}$ 5d SCFT, then $S$ is the corresponding del Pezzo surface obtained from $\mathbb{P}^{2}$ by blowing up $Q-3$ points in general position. On the other hand, if $(Y,D,L)$ corresponds to the $\widetilde{E}_{1}$ 5d SCFT, then $S$ is the corresponding del Pezzo surface $\mathbb{F}_{0}=\mathbb{P}^{1}\times\mathbb{P}^{1}$. In particular, we recover the description of $E_{n}$ and $\widetilde{E}_{n}$ 5d SCFTs as M-theory on the canonical 3-fold singularity obtained by contracting to a point the zero-section of a local del Pezzo surface [83]. If $k>1$, then $\mathcal{X}$ is a crepant resolution of the order $k$ base change $t\mapsto t^{k}$ of $K_{S}\rightarrow\Delta$. The base change creates a family of $A_{k-1}$-singularities parametrized by $E$, whose resolution produces $k-1$ additional irreducible components in $\mathcal{X}_{0}$, all isomorphic to $\mathbb{P}^{1}$-bundles over $E$.
@@ 6.3. Calabi–Yau 3-folds and smoothings of slc singularities
[…]
Figure 6.3. The slc surface $\mathcal{X}_{0}^{\mathrm{can}}$ and its minimal resolution: degenerate cusp on the left, cusp in the middle, and simple elliptic singularity on the right.
Let $(Y,D,L)$ be the log Calabi–Yau surface associated to $W^{\mathrm{asym}}$. If $L^{2}>0$ and $L\cdot D=0$, that is $W^{\mathrm{asym}}\in\mathbf{CWebs}_{I}$, then we can consider the corresponding polarized log Calabi–Yau surface $(Y^{\mathrm{pol}},D^{\mathrm{pol}},L^{\mathrm{pol}})$ as in Equation (4.8). The 0-dimensional strata $x_{i}$ of $D^{\mathrm{pol}}$ are in one-to-one correspondence with the vertices $v_{i}$ of a Symington polytope $P$ attached to $W^{\mathrm{asym}}$. Moreover, each 0-dimensional strata admits a neighborhood isomorphic to an affine toric surface $Y_{i}$, with fan given by the dual cone of the cone spanned by the edges of $P$ at the vertex $v_{i}$. Let $X_{i}$ be the dual affine toric surface, with fan given by the cone spanned by the edges of $P$ at the vertex $v_{i}$. Then, we have
@@
that is, the irreducible components of the surface $\mathcal{X}_{0}^{\mathrm{can}}$ are naturally indexed by the vertices of $P$, and are isomorphic to the toric surfaces $X_{i}$. Moreover, these surfaces are glued together along their toric divisors according to the cyclic order defined by the boundary of $P$, and all intersect together at the point $x_{0}$, see Figure 6.3. Each irreducible component $X_{i}$ is a quotient cyclic singularity, and so in this case $\mathcal{X}_{0}^{\mathrm{can}}$ is a degenerate cusp surface singularity in the sense of [91, §1].
If $L^{2}>0$ and $L\cdot D=0$, that is $W^{\mathrm{asym}}\in\mathbf{CWebs}_{II,+}$, then we can consider the corresponding Calabi–Yau surface $(Y^{\mathrm{pol}},L^{\mathrm{pol}})$ with a cusp singularity obtained by contracting the cycle of rational curves $D$ to a point as in Equation (4.9). Recall that a cusp singularity is a normal surface singularity whose exceptional divisor in the minimal resolution is a cycle of rational curves. Cusp singularities come in dual pairs [84] and we refer to [76, §2], [47, §7.1], [36, §2] for the explicit description of this duality. In this case, the surface $\mathcal{X}_{0}^{\mathrm{can}}$ is normal, with an isolated singularity at the point $x_{0}$, which is the cusp surface singularity dual to the cusp surface singularity of $Y^{\mathrm{pol}}$.
@@
Degenerate cusp, cusp and simple elliptic singularities are examples of slc (semi-log-canonical) surface singularities as introduced in [70, Definition 4.17]. The notion of slc surface singularity is closely related to the notion of canonical 3-fold singularity. Indeed, if $\mathcal{Z}_{0}$ is surface singularity with a smoothing $\mathcal{Z}\rightarrow\Delta$ that admits a semi-stable resolution, then, by Theorem [70, Theorem 5.1], the surface $\mathcal{Z}_{0}$ is slc if and only if the 3-fold $\mathcal{Z}$ has canonical singularities. In other words, slc surface singularities are the surface singularities whose one-parameter smoothings produce canonical 3-fold singularities. According to [70, Theorem 4.21], degenerate cusp, cusp, and simple elliptic singularities are the three main classes of Gorenstein slc singularities, that is, with an invertible canonical sheaf. General slc singularities are only $\mathbb{Q}$-Gorenstein in general, that is, only a positive power of their canonical sheaf is a line bundle. The non-Gorenstein slc singularities are classified in [70, Theorem 4.24]: they are either $\mathbb{Z}/2\mathbb{Z}$, $\mathbb{Z}/3\mathbb{Z}$, $\mathbb{Z}/4\mathbb{Z}$ or $\mathbb{Z}/6\mathbb{Z}$ quotients of simple elliptic singularities, or $\mathbb{Z}/2\mathbb{Z}$ quotients of cusp and degenerate cusp singularities. We expect the canonical 3-fold singularities defined by one-parameter smoothings of these non-Gorentsein slc singularities to be M-theory duals of webs of 5-branes with 7-branes and additional orientifolds and S-folds, as in [20, 98] and [1, 66] respectively, but we leave the study of this correspondence to future work.

### v2

@@ 6.2. Calabi–Yau 3-folds as non-toric deformations of toric Calabi–Yau 3-folds
in $\mathbb{C}^{3}$ by the group $\mathbb{Z}/n\mathbb{Z}$ acting with weights (1,1,a)(1,-1,a) on $\mathbb{C}^{3}$. It follows from this description that $(\overline{X}_{\mathcal{C}},\partial{\overline{X}}_{\mathcal{C}})$ admits a natural one-parameter deformation given by quotienting the hypersurface
@@
by the action of the group $\mathbb{Z}/n\mathbb{Z}$ acting with weights (1,1,a)(1,-1,a) on $\mathbb{C}^{3}$. For $n=1$, this is just smoothing the normal crossing divisor $xy=0$ into $xy=t$. If $n>1$, then, both the singular toric surface $\overline{X}_{\mathcal{C}}$ and its toric boundary divisor $\partial\overline{X}_{\mathcal{C}}$ become smooth for $t\neq 0$.
According to [6, Lemma 4.1], there is no obstruction to applying this deformation simultaneously to all neighborhoods $(\overline{X}_{\mathcal{C}},\partial{\overline{X}}_{\mathcal{C}})$ of the 0-dimensional strata $x_{\mathcal{C}}$. This yields a global deformation of the toric surface $(\overline{X}_{0}^{v},\partial\overline{X}_{0}^{v})$, and we denote the resulting log Calabi–Yau surface by $(X_{0}^{v},\partial X_{0}^{v})$. This surface is smooth since all the possibly singular 0-dimensional strata of $(\overline{X}_{0}^{v},\partial\overline{X}_{0}^{v})$ become smooth in the deformation. Here, we adopt a slightly broader interpretation of “log Calabi–Yau” surface than in §3.1.2. First, the surface $(\overline{X}_{0}^{v},\partial\overline{X}_{0}^{v})$ may be non-compact, as occurs when $v$ lies on the boundary of $P$. Moreover, the boundary divisor $\partial\overline{X}_{0}^{v}$ may be smooth. which occurs precisely when all cones in the star of $v$ have been cut out; this situation arises exactly when $L^{2}=0$ – see Example 7.4.
In the web picture, the vertex $v$ of $\mathcal{P}$ corresponds to a face of the polyhedral decomposition defined by $W^{\prime}$. This face arises from a corresponding face of the polyhedral decomposition defined by $W$ through a series of “corner smoothings” induced by pushing-in 7-branes, as illustrated in Figure 22.
Figure 22. The log Calabi–Yau surface $(X_{0}^{v},\partial X_{0}^{v})$ obtained by “smoothing corners” of the toric surface $(\overline{X}_{0}^{v},\partial\overline{X}_{0}^{v})$.
When not all cones in the star of $v$ are cut out, the log Calabi–Yau surface $(X_{0}^{v},\partial X_{0}^{v})$ admits a more explicit description. The singularity of the integral affine manifold $P$ at $v$ can be viewed as several focus-focus singularities brought together, one for each cut triangle with tip at $v$. Since not all cones of the star of $v$ are cut out, these focus-focus singularities can be slightly displaced along their monodromy invariant directions, effectively smoothing the integral affine structure at $v$. After this smoothing, the star of $\mathcal{P}$ at $v$ defines the fan of a toric variety. The log Calabi-Yau surface $(X_{0}^{v},\partial X_{0}^{v})$ is then obtained from this toric variety as follows:
@@
    First, perform toric blow-ups so that every half-line emanating from $v$ and passing through a focus-focus singularity is incorporated into the fan.
@@
    Then, for each focus-focus singularity, perform a non-toric blow-up at a point on the toric boundary component corresponding to the relevant half-line.
@@
    Finally, contract all boundary components that do not correspond to edges of $\mathcal{P}$.
In general, there are multiple ways to decompose the integral affine singularity at $v$ into focus-focus singularities, corresponding to distinct toric models of the log Calabi–Yau surface $(X_{0}^{v},\partial X_{0}^{v})$.
When all cones in the star of $v$ are cut out, then we have $L^{2}=0$. As shown in Example 6.1, the resulting surface $X_{0}^{v}$ is a smooth del Pezzo surface, and $\partial X_{0}^{v}$ is a smooth anticanonical divisor in $X_{0}^{v}$ – see also Example 7.4.
Step II: Construction of the normal crossing surface $\mathcal{X}_{0}$. By Step I, a smooth log Calabi–Yau surface $(X_{0}^{v},\partial X_{0}^{v})$ has been assigned to every vertex of $\mathcal{P}$, with dual intersection given by the star of $\mathcal{P}$ at $v$. In particular, the edges of $\mathcal{P}$ adjacent to $v$ correspond to irreducible components of the boundary divisor $\partial X_{0}^{v}$. Gluing the surfaces $(X_{0}^{v},\partial X_{0}^{v})$ along boundary components associated with the same edges of $\mathcal{P}$ yields a singular surface $\mathcal{X}_{0}$, whose irreducible components are the various $(X_{0}^{v},\partial X_{0}^{v})$ and whose intersection complex is $\mathcal{P}$. Equivalently, the intersection complex of $\mathcal{X}_{0}$ coincides with the polyhedral decomposition determined by the web $W^{\prime}$. Since each face of of $\mathcal{P}$ is a triangle, reflecting the 3-valency of $W^{\prime}$, there are at most triple intersections among the irreducible components of $(X_{0}^{v},\partial X_{0}^{v})$. As each $X_{0}^{v}$ is smooth, it follows that $\mathcal{X}_{0}$ is a normal crossing surface.
Step III: Construction of the Calabi–Yau 3-fold $\mathcal{X}$. We obtain a smooth non-compact Calabi–Yau 3-fold $\mathcal{X}$ as the total space of a one-parameter smoothing of the normal crossing surface. Let $(Y,D,L)$ be the log Calabi–Yau surface associated to the asymptotic web of 5-branes with 7-branes $W^{\mathrm{asym}}$. As established in [39] for the case $L\cdot D=0$, and in [6, Theorem 3.36] for $L\cdot D>0$, there exists a smooth non-compact Calabi–Yau 3-fold $\mathcal{X}$ and a holomorphic map
@@
where $\Delta$ is an open disk in $\mathbb{C}$ containing $0\in\mathbb{C}$, such that the central fiber $\pi^{-1}(0)$ is a reduced normal crossing divisor isomorphic to $\mathcal{X}_{0}$. Furthermore, it is shown in [40, Proposition 3.14] for the case $L\cdot D=0$, and in [7] for $L\cdot D>0$, that for $t\neq 0$, the general fiber $\pi^{-1}(t)$ is diffeomorphic to $U=Y\setminus D$. Finally, by [6, Theorem 3.29], there exists a unique holomorphic map
@@
such that $\pi=\pi^{\mathrm{can}}\circ f$. Moreover, $\mathcal{X}^{\mathrm{can}}$ is non-compact Calabi–Yau 3-fold with Gorenstein canonical singularities, and $f:\mathcal{X}\rightarrow\mathcal{X}^{\mathrm{can}}$ is a crepant resolution. By construction, $\mathcal{X}^{\mathrm{can}}$ is a generally non-toric deformation of the toric 3-fold singularity $\overline{\mathcal{X}}^{\mathrm{can}}$ associated as in §2.3 to the asymptotic web of 5-branes $\overline{W}^{\mathrm{can}}$. Moreover, the crepant resolution $f:\mathcal{X}\rightarrow\overline{X}^{\mathrm{can}}$ is a deformation of the crepant toric partial resolution $\overline{\mathcal{X}}\rightarrow\overline{\mathcal{X}}^{\mathrm{can}}$ associated to the web of 5-branes $\overline{W}$. The toric Calabi–Yau 3-fold $\overline{\mathcal{X}}$ is singular in general since $\overline{W}$ might contain 3-valent vertices of multiplicity >1>1 or 4-valent vertices. Thus, $\mathcal{X}$ is a smoothing of $\overline{\mathcal{X}}$.
@@
Assume that $L^{2}=0$, that is, as reviewed above Equation (4.10) that there exists an elliptic fibration $c:Y\rightarrow\mathbb{P}^{1}$ with $D=c^{-1}(\infty)$ and $L=\mathcal{O}(kD)$. In particular a general curve $C\in|L|$ consists of $k$ disjoint elliptic fibers. In this case, the web $W^{\prime}$ obtained after fully pushing the 7-branes does not contain any vertex, and is topologically the union of $k$ concentric circles, which can be viewed as the tropicalization of the $k$ elliptic connected components of $C$. If $k=1$, then the complement of $W^{\prime}$ contains a unique bounded region, corresponding to a compact component of $\mathcal{X}_{0}$, which is necessarily a del Pezzo surface $S$. It follows that $\mathcal{X}$ is the total space $K_{S}$ of the canonical divisor of $S$, that is, a “local del Pezzos surface”, and the map $\pi:\mathcal{X}\rightarrow\Delta$ is induced by a general anticanonical section, whose zero-locus is a smooth genus one curve $E$. If $(Y,D,L)$ corresponds to the $E_{Q-3}$ 5d SCFT, then $S$ is the corresponding del Pezzo surface obtained from $\mathbb{P}^{2}$ by blowing up Q3Q-3 points in general position. On the other hand, if $(Y,D,L)$ corresponds to the $\widetilde{E}_{1}$ 5d SCFT, then $S$ is the corresponding del Pezzo surface $\mathbb{F}_{0}=\mathbb{P}^{1}\times\mathbb{P}^{1}$. In particular, we recover the description of $E_{n}$ and $\widetilde{E}_{n}$ 5d SCFTs as M-theory on the canonical 3-fold singularity obtained by contracting to a point the zero-section of a local del Pezzo surface [86]. If $k>1$, then $\mathcal{X}$ is a crepant resolution of the order $k$ base change $t\mapsto t^{k}$ of $K_{S}\rightarrow\Delta$. The base change creates a family of $A_{k-1}$-singularities parametrized by $E$, whose resolution produces k1k-1 additional irreducible components in $\mathcal{X}_{0}$, all isomorphic to $\mathbb{P}^{1}$-bundles over $E$.
@@ 6.3. Calabi–Yau 3-folds and smoothings of slc singularities
Figure 23. The slc surface $\mathcal{X}_{0}^{\mathrm{can}}$ and its minimal resolution: degenerate cusp on the left, cusp in the middle, and simple elliptic singularity on the right.
Let $(Y,D,L)$ be the log Calabi–Yau surface associated to $W^{\mathrm{asym}}$. If $L^{2}>0$ and $L\cdot D>0$, that is $W^{\mathrm{asym}}\in\mathbf{CWebs}_{I}$, then we can consider the corresponding polarized log Calabi–Yau surface $(Y^{\mathrm{pol}},D^{\mathrm{pol}},L^{\mathrm{pol}})$ as in Equation (4.8). The 0-dimensional strata $x_{i}$ of $D^{\mathrm{pol}}$ are in one-to-one correspondence with the vertices $v_{i}$ of a Symington polytope $P$ attached to $W^{\mathrm{asym}}$. Moreover, each 0-dimensional strata admits a neighborhood isomorphic to an affine toric surface $Y_{i}$, with fan given by the dual cone of the cone spanned by the edges of $P$ at the vertex $v_{i}$. Let $X_{i}$ be the dual affine toric surface, with fan given by the cone spanned by the edges of $P$ at the vertex $v_{i}$. Then, we have
@@
that is, the irreducible components of the surface $\mathcal{X}_{0}^{\mathrm{can}}$ are naturally indexed by the vertices of $P$, and are isomorphic to the toric surfaces $X_{i}$. Moreover, these surfaces are glued together along their toric divisors according to the cyclic order defined by the boundary of $P$, and all intersect together at the point $x_{0}$, see Figure 23. Each irreducible component $X_{i}$ is a quotient cyclic singularity, and so in this case $\mathcal{X}_{0}^{\mathrm{can}}$ is a degenerate cusp surface singularity in the sense of [94, §1].
If $L^{2}>0$ and $L\cdot D=0$, that is $W^{\mathrm{asym}}\in\mathbf{CWebs}_{II,+}$, then we can consider the corresponding Calabi–Yau surface $(Y^{\mathrm{pol}},L^{\mathrm{pol}})$ with a cusp singularity obtained by contracting the cycle of rational curves $D$ to a point as in Equation (4.9). Recall that a cusp singularity is a normal surface singularity whose exceptional divisor in the minimal resolution is a cycle of rational curves. Cusp singularities come in dual pairs [87] and we refer to [79, §2], [50, §7.1], [39, §2] for the explicit description of this duality. In this case, the surface $\mathcal{X}_{0}^{\mathrm{can}}$ is normal, with an isolated singularity at the point $x_{0}$, which is the cusp surface singularity dual to the cusp surface singularity of $Y^{\mathrm{pol}}$.
@@
Degenerate cusp, cusp and simple elliptic singularities are examples of slc (semi-log-canonical) surface singularities as introduced in [73, Definition 4.17]. The notion of slc surface singularity is closely related to the notion of canonical 3-fold singularity. Indeed, if $\mathcal{Z}_{0}$ is surface singularity with a smoothing $\mathcal{Z}\rightarrow\Delta$ that admits a semi-stable resolution, then, by Theorem [73, Theorem 5.1], the surface $\mathcal{Z}_{0}$ is slc if and only if the 3-fold $\mathcal{Z}$ has canonical singularities. In other words, slc surface singularities are the surface singularities whose one-parameter smoothings produce canonical 3-fold singularities. According to [73, Theorem 4.21], degenerate cusp, cusp, and simple elliptic singularities are the three main classes of Gorenstein slc singularities, that is, with an invertible canonical sheaf. General slc singularities are only $\mathbb{Q}$-Gorenstein in general, that is, only a positive power of their canonical sheaf is a line bundle. The non-Gorenstein slc singularities are classified in [73, Theorem 4.24]: they are either $\mathbb{Z}/2\mathbb{Z}$, $\mathbb{Z}/3\mathbb{Z}$, $\mathbb{Z}/4\mathbb{Z}$ or $\mathbb{Z}/6\mathbb{Z}$ quotients of simple elliptic singularities, or $\mathbb{Z}/2\mathbb{Z}$ quotients of cusp and degenerate cusp singularities. We expect the canonical 3-fold singularities defined by one-parameter smoothings of these non-Gorentsein slc singularities to be M-theory duals of webs of 5-branes with 7-branes and additional orientifolds and S-folds, as in [23, 101] and [1, 69] respectively, but we leave the study of this correspondence to future work.